There are two triples of positive integers (a,b,c and d,e,f) such that
a²+b²+c²=86 and d²+e²+f²=86. Evaluate the expression absolute value of
abc-def.

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

There are two triples of positive integers (a,b,c and d,e,f) such that
a²+b²+c²=86 and d²+e²+f²=86. Evaluate the expression absolute value of
------------------
Obviously, all are less than 10.
There's no algebraic solution, just trial and error.
There's no reason for me to do it.
You do it.

Answer by ikleyn(52777)

.
1)  86 = 1^2 + 2^2 + 9^2


2)  86 = 5^2 + 5^2 + 6^2.


You do the rest.